OpenStudy (anonymous):

the legs of an isosceles triangle are 13 units long and the altitude to the base is 12 units long. find the length of the base.

OpenStudy (anonymous):

first draw it and you'll notice that the altitute bisects the triangle into 2 equal triangles so the bases will be equal, and so, use the pythagoreous theorem to find the base of one and the other one will be equal to it, then add, so you'll have this :\[c^2 = y^2 + x^2\] we're solving for x, so : \[x^2 = c^2 -y^2 = \sqrt{13^2 - 12^2}=\sqrt{25}= 5\] we have found the first base = 5, the other base is equal to it so : Base = B1 + B2 = 5 + 5 = 10 units ^_^

OpenStudy (anonymous):

that makes more sence thanks

OpenStudy (anonymous):

np :)